Calculate electric field intensity which would just sufficient to balance the weight of the particle of charge-10micro Coulomb and mass10mg.

Grade:12

2 Answers

Ananth Narayanan
7 years ago
For the given condition m*g=k*q*E where k=9*10^9.
Therefore E=(m*g)/k*q
                 =(10^-2 *9.8)/(9* 10^9 * 10^-5)
                 =1.088 *10^-6
Abhishek
5 years ago
Calculate the electric field intensity which could be just sufficient to balance the weight of an electron